OpenStudy (anonymous):

Two protons are released from rest when they are 0.750 nm apart. What is the maximum speed they will reach?

OpenStudy (irishboy123):

i would compare the energies at r = r_o and r = \(\infty\) \(k\frac{q.q}{r_o} = 2(\frac{1}{2}m\dot r_{\infty}^2)\)

OpenStudy (irishboy123):

http://openstudy.com/study#/updates/55fe508fe4b0ed58e2759c0f this is also being pursued in physics